# Mitigating distribution shift in machine learning-augmented hybrid simulation

### v1

@@ 4 Numerical experiments
In this section, we present numerical experiments which illustrate the effect of the tangent-space regularized algorithm presented in Algorithm 1. We use the experiments to illustrate two phenomena. The first demonstrates the effectiveness of our algorithm, for which we test on the reaction-diffusion equation and the incompressible Navier-Stokes equation, and compare with benchmark algorithms using least squares, as well as simple regularizations that do not account for the structure of the resolved dynamics or the data manifold. For the second goal, we show that our method brings more significant improvements in problems with severe distribution shifts. This is demonstrated on the same family of PDEs, i.e. reaction-diffusion equation with varying diffusion coefficient and Navier-Stokes equation with different Reynolds numbers. These varying parameters can be understood - as we show numerically - as quantitative indicators of the severity of distribution shift. The implementation of our method and experiment reproduction is found in the repository [57].
@@ 4.1 Tangent-space regularized algorithm improves simulation accuracy
We first show that the proposed algorithm can be advantageous on several prototypical machine learning augmented scientific computing applications.
In the first experiment, we use the FitzHugh-Nagumo reaction-diffusion Eq. 11 equation as the test case of our algorithm. The resolved-unresolved structure is the coarse-fine grid correction structure discussed in Section 2.2. This helps us obtain high-fidelity simulation results by combining a low-fidelity numerical solver with a fast neural network surrogate model for correction. Here we choose $\alpha=0.01,\beta=1.0,D=\begin{pmatrix}\gamma&0\\ 0&2\gamma\end{pmatrix},\gamma\in\{0.05,0.10,0.15,0.20,0.25\}$ in the simulation which guarantee the existence of Turing patterns. The initial condition $\mathbf{u}_{0}$ is generated by i.i.d. sampling from a normal distribution:
$u(x,y,0),v(x,y,0)\sim\mathcal{N}(0,1).$ (34)
This system is known to have a dynamical spatial patterns where initial smaller “clusters” gradually transforms and merges to form bigger cluster according to the balance between diffusion and reaction effects [33]. In literature, the importance of the sampling procedure is also emphasized and various adaptive sampling method based on error and particle methods are studied in [12, 52, 58].
@@
where the mapping $\phi$ between the fine grid state $\mathbf{u}_{k}^{2n}$ to the correction term $\mathbf{y}_{k}^{2n}$ is the unresolved model we desire to learn. Here $I_{n}^{2n},R_{2n}^{n}$ are fixed restriction and interpolation operators between two meshes, i.e.
$R_{2n}^{n}:\mathbb{R}^{2n\times 2n}\rightarrow\mathbb{R}^{n\times n},\quad\mathbf{u}_{ij}^{n}=\frac{\mathbf{u}_{2i,2j}^{2n}+\mathbf{u}_{2i,2j+1}^{2n}+\mathbf{u}_{2i+1,2j}^{2n}+\mathbf{u}_{2i+1,2j+1}^{2n}}{4},$ (36)
@@
To obtain training data for $\mathbf{u}_{n}^{t},\mathbf{u}_{2n}^{t}$, we simulate Eq. 11 (with the same initial condition) on two meshes of sizes $64\times 64$ and $128\times 128$ respectively, with time step $\Delta t=0.01,N_{\text{step}}=T/\Delta t=100$, using a second order Crank-Nicolson scheme. It remains to calculate the labels for the training data, i.e. $\mathbf{y}_{k}^{2n}=\phi(\mathbf{u}_{k}^{2n})$. These are obtained by using the first half of the Eq. 36, i.e.
$\mathbf{y}_{n}^{t}=\mathbf{u}_{2n}^{t+1}-I_{n}^{2n}\circ f_{n}(R_{2n}^{n}(\mathbf{u}_{2n}^{t}))),$ (37)
@@
Finally, we pair them together to form the training data $\left\{(\mathbf{u}_{n}^{1},\mathbf{y}_{n}^{1}),\cdots,(\mathbf{u}_{n}^{N},\mathbf{y}_{n}^{N})\right\}$. The data here is different from the one introduced in Section 2.3 since the system is autonomous, so we do not include time indices. To summarize, the training data contains two patches of field component patterns of size $100\times 128\times 128$ corresponding to $\mathbf{u},\mathbf{v}$ respectively and two patches $\mathbf{y}_{u},\mathbf{y}_{v}$ of the same shape and generated according to Eq. 36.
@@
The baseline algorithms we implemented to compare with ours is $\phi_{\text{OLS}}$, which minimizes the mean-squared error on the training data, i.e.
$\phi_{\text{OLS}}=\arg\min\mathbb{E}_{\mathbf{u}\sim\rho_{0}}\left\|\phi_{\theta}(\mathbf{u})-\phi(\mathbf{u})\right\|_{2}^{2},$ (38)
@@
where we use $\rho_{0}$ to denote the distribution of the field variable generated by certain simulation condition Eq. 34. The corresponding numerical results are labelled by “OLS”.
@@
Figure 3: Simulation of the reaction-diffusion equation and configurations at time step: 0, 100, 200, 300, 500, (a) ground truth; (b) simulated fluid field using OLS estimator; (c) simulated fluid field using TR estimator. (d) is the comparison of the ordinary least squares and regularized one on reaction-diffusion equation: the solid line represents the averaged error along the trajectory and the dash line represented the averaged distribution shift calculated using Eq. 24. The shadow part is the standard deviation in ten random experiments with different training and testing data. Both the error and the distribution shift increase at a lower speed in TR simulation than in OLS simulation. This suggests that the better performance of TR algorithm in reaction-diffusion equation is partially caused by improving the distribution shift issue.
@@
We summarize the comparison results in Fig. 3. Here, solid lines represent the errors along the trajectories and dashed lines represent the distribution shift. The shaded part is the standard deviation over ten random experiments with different split of the training and testing data. As can be observed, compared to our method the OLS baseline incur greater errors and distribution shifts. Moreover, this difference increases with time. In contrast, consistent with our analysis in the simplified setting, our algorithm mitigates both the distribution shift and the trajectory error.
@@
Next, we consider the Navier-Stokes equation Eq. 3 as the test case of our algorithm Algorithm 1. In this case, the resolved model is an explicit scheme to evolve the velocity field while unresolved model is calculated via a fast CNN surrogate, which can accelerate the simulation for large grid system. The simulation domain is rectangular with aspect ratio $1/4$. On the upper and lower boundary we impose no-slip boundary condition on the velocity and on the inlet, i.e. $x=0$ we specify the velocity field $\mathbf{u}$. Moreover, at the outlet, i.e. $x=n$ we pose zero-gradient conditions on both the horizontal velocity and pressure, i.e. $\frac{\partial p}{\partial n}=\frac{\partial u}{\partial n}=0$. The inflow boundary is defined according to following functional forms:
$\mathbf{u}(0,y,t)=\begin{pmatrix}u(0,y,t)\\ v(0,y,y)\end{pmatrix}=\begin{pmatrix}\exp\{-50(y-y_{0})^{2}\}\\ \sin t\cdot\exp\{-50(y-y_{0})^{2}\}\end{pmatrix}$ (39)
@@
The training configuration is calculated using the projection method [10] with staggered grid, i.e. the pressure is placed at the center of each cell and horizontal (vertical) velocity is placed at vertical (horizontal) edge of each cell. The jet location $y_{0}$ is selected from $0.3$ to $0.7$ uniformly.
@@
Again, we compare the performance of our algorithm and the OLS method. As illustrated in Fig. 4, naively using OLS to estimate the unresolved models will generally lead to error blow-up, whereas using our regularized estimator we can obtain a reasonable simulation along the whole time period. Moreover, the trend of the trajectory error fits that of the distribution shift. In particular, the error and distribution shift blow-ups occur simultaneously. Our algorithm again decreases the trajectory error by mitigating the problem of distribution shift.
@@
Figure 4: Simulation of the Navier-Stokes equation and configurations at time step: 0, 60, 120, 180, 240, (a) ground truth; (b) simulated fluid velocity field using OLS estimator; (c) simulated fluid velocity field using TR estimator; (d) the comparison of the OLS estimator and the regularized one. The solid lines represent the trajectory error and the dashed lines represent distribution shift calculated using Eq. 24. We observe that the OLS estimator leads to error (and DS) blow-up at around 270 time steps, while our method remains stable. Again, the trends of the error and the distribution shift are highly correlated, and our method that controls distribution shift leads to error control, consistent with our analyses.
@@ 4.2 Performance under distribution shift of different magnitudes
We now demonstrate concretely that our method brings more significant improvements when the intrinsic distribution shift is more severe. Intuitively speaking, since our algorithm penalizes the deviation of the simulated trajectories to the data manifold while OLS totally ignores this, it would be reasonable to expect that our algorithm will outperform OLS more in the problems where such deviations are large. A sanity check showing different distribution shift extents under different parameters is provided in the supplementary materials LABEL:fig:NSRe-ds.
Here, we introduce two more baselines for ablation studies. Comparing with the OLS algorithm, these algorithms both add some regularization to the loss function while their regularizations are general and not dynamics-specific compared to our regularization. Consequently, any improved performance in our algorithm over these regularized algorithms suggests that targeted regularization, which takes into account the structure of the resolved part, effectively balances accuracy and stability.
The first baseline regularizes the original least squares objective by a term which quantifies some complexity the network model, e.g.
$\phi_{\text{mOLS}}=\arg\min\mathbb{E}_{\mathbf{u}\sim\rho_{0}}\left\|\phi_{\theta}(\mathbf{u})-\phi(\mathbf{u})\right\|_{2}^{2}+\lambda\left\|\theta\right\|_{2}^{2}.$ (40)
Here the naive $L^{2}$ regularized algorithm is implemented by adjusting the weight decay [31] parameter of the Adam solver during the training process. The corresponding numerical results are labelled by “mOLS” (“m” for modified).
@@
$\phi_{\text{aOLS}}=\arg\min\mathbb{E}_{\mathbf{u}\sim\rho_{\xi}}\left\|\phi_{\theta}(\mathbf{u})-\phi(\mathbf{u})\right\|_{2}^{2}.$ (41)
Notice that the training data follows the distribution $\rho_{\epsilon}=\rho_{0}*\xi$ where *** denotes the convolution operator and $\xi$ is some random noise which we take to be Gaussian random variable with an adjustable variance. In implementation, we generate the noise and add to the input during each epoch so the inputs are not the same across epochs. Moreover, this noise $\xi$ should be distinguished from the observational noise $\epsilon$ we introduced in Section 3. This method is used in several previous work [46, 36] to improve stability and prevent the error from accumulating too fast. As this technique is similar to adversarial attack and data augmentation [44, 14, 13], we name this estimator and algorithm as “aOLS” (“a” stands for adversarial). We slightly abuse the terminology here by naming this random perturbation as some kind of adversarial attack.
Table 1: Performance on reaction-diffusion equation over 10 experiments.
@@
In Table 1, we test the algorithms under different diffusion coefficient $\gamma=0.05$, $0.10,0.15,0.20,0.25$. We calculate the relative trajectory error at time $T=1000$, i.e. $\frac{\left\|\mathbf{u}_{k}-\widehat{\mathbf{u}}_{k}\right\|}{\left\|\mathbf{u}_{k}\right\|}$ over ten random split of the training and test data. The last row calculates the performance difference between our method and the OLS baseline. We observe that under almost all the settings our method out-performs OLS, mOLS, aOLS baselines, and comparing horizontally, we see that the relative improvement decreases with $\gamma$, as diffusion term mitigate the distribution shift.
We now consider the Navier-Stokes equation. Since the baseline method quickly leads to error blow up, we use another comparison criterion. We define a stopping time $t_{K}=\arg\max_{t}\frac{\left\|\mathbf{u}_{t}-\widehat{\mathbf{u}}_{t}\right\|}{\left\|\mathbf{u}_{t}\right\|}\leq K$ where $K$ is a error threshold to be determined during the experiments. We calculate the first time the trajectory error reaches a threshold under different Reynolds numbers and different mesh sizes. The results are shown in Table 2. Under almost all the scenarios, our method outperforms all the baseline algorithms by a larger $t_{K}$. This validates the effectiveness of our algorithm. Comparing Table 2 horizontally, one finds that the improvement of our method is also increasing with respect to the Reynolds number. As the flow field generally becomes more complex (and sensitive) with increasing Reynolds number, our method is expected to bring bigger improvements. This is consistent with our experiments.
@@
TR $\mathbf{506\pm 18}$ $\mathbf{497\pm 28}$ $\mathbf{482\pm 29}$ $\mathbf{442\pm 21}$ $\mathbf{452\pm 13}$
@@
TR $636\pm 35$ $\mathbf{602\pm 28}$ $585\pm 23$ $\mathbf{422\pm 35}$ $\mathbf{402\pm 28}$

### v2

@@ 4 Numerical experiments
In this section, we present numerical experiments which illustrate the effect of the tangent-space regularized algorithm presented in Algorithm 1. We use the experiments to illustrate two phenomena. The first demonstrates the effectiveness of our algorithm, which we test over reaction-diffusion equations and the incompressible Navier-Stokes equation, and compare with benchmark algorithms using the least squares, as well as simple regularizations that do not account for the structure of the resolved dynamics or the data manifold. For the second goal, we show that our method brings more significant improvements in problems with severe distribution shifts. This is demonstrated on the same family of PDEs, i.e. reaction-diffusion equations with varying diffusion coefficients and the Navier-Stokes equation with different Reynolds numbers. These varying parameters can be understood - as we show numerically - as quantitative indicators of the severity of distribution shift. The implementation of our method and experiment reproduction is found in the repository [65].
@@ 4.1 Tangent-space regularized algorithm improves simulation accuracy
We first show that the proposed algorithm can be advantageous in several prototypical machine-learning augmented scientific computing applications.
In the first experiment, we use the FitzHugh-Nagumo reaction-diffusion equation Eq. 11 as the test case of our algorithm. The resolved-unresolved structure is the coarse-fine grid correction structure discussed in Section 2.2. This helps us obtain high-fidelity simulation results by combining a low-fidelity numerical solver with a fast neural network surrogate model for correction. Here we choose $\alpha=0.01,\beta=1.0,D=\begin{pmatrix}\gamma&0\\ 0&2\gamma\end{pmatrix},\gamma\in\{0.05,0.10,0.15,0.20,0.25\}$ in the simulation which guarantees the existence of Turing patterns. The initial condition $\mathbf{u}_{0}$ is generated by i.i.d. sampling from a normal distribution $u(x,y,0),v(x,y,0)\sim\mathcal{N}(0,1)$. This system is known to have dynamic spatial patterns where initial smaller clusters gradually transform and merge to form bigger clusters according to the balance between diffusion and reaction effects [36]. In literature, the importance of the sampling procedure is also emphasized, and various adaptive sampling methods based on error and particle methods are studied in [12, 60, 66].
@@
The mapping $\phi$ between the fine grid state $\mathbf{u}_{k}^{2n}$ to the correction term $\mathbf{y}_{k}^{2n}$ is the unresolved model we desire to learn. Here $I_{n}^{2n},R_{2n}^{n}$ are fixed restriction and interpolation operators between two meshes, i.e.
(48) $R_{2n}^{n}:\mathbb{R}^{2n\times 2n}\rightarrow\mathbb{R}^{n\times n},\quad\mathbf{u}_{ij}^{n}=\frac{\mathbf{u}_{2i,2j}^{2n}+\mathbf{u}_{2i,2j+1}^{2n}+\mathbf{u}_{2i+1,2j}^{2n}+\mathbf{u}_{2i+1,2j+1}^{2n}}{4},$
@@
To obtain training data for $\mathbf{u}_{n}^{t},\mathbf{u}_{2n}^{t}$, we simulate Eq. 11 (with the same initial condition) on two meshes of sizes $64\times 64$ and $128\times 128$ respectively, with time step $\Delta t=0.01,N_{\text{step}}=T/\Delta t=100$, using a second-order Crank-Nicolson scheme. It remains to calculate the labels for the training data, i.e. $\mathbf{y}_{k}^{2n}=\phi(\mathbf{u}_{k}^{2n})$. These are obtained by using the first half of the Eq. 48. Finally, we pair them together to form the training data $\left\{(\mathbf{u}_{n}^{1},\mathbf{y}_{n}^{1}),\cdots,(\mathbf{u}_{n}^{N},\mathbf{y}_{n}^{N})\right\}$. The data here is different from the one introduced in Section 2.3 since the system is autonomous, so we do not include time indices. To summarize, the training data contains two patches of field component patterns of size $100\times 128\times 128$ corresponding to $\mathbf{u},\mathbf{v}$ respectively and two patches $\mathbf{y}_{u},\mathbf{y}_{v}$ of the same shape and generated according to Eq. 48.
@@
The baseline algorithms we implemented to compare with ours is $\phi_{\mathrm{OLS}}$, which minimizes the mean-squared error on the training data, i.e.
(49) $\phi_{\mathrm{OLS}}=\arg\min\mathbb{E}_{\mathbf{u}\sim\rho_{0}}\left\|\phi_{\theta}(\mathbf{u})-\phi(\mathbf{u})\right\|_{2}^{2},$
@@
where we use $\rho_{0}$ to denote the distribution of the field variable. The corresponding numerical results are labeled by “OLS”.
@@
Figure 4: Simulation of reaction-diffusion equations and configurations at time step: 0, 100, 200, 300, 500, (a) ground truth; (b) simulated fluid field using OLS estimator; (c) simulated fluid field using $\mathrm{TR}$ estimator. (d) is the comparison of the ordinary least squares and regularized one on reaction-diffusion equations: the solid line represents the averaged error along the trajectory and the dashed line represents the averaged distribution shift calculated using function $F$ defined in Eq. 20. The shadow part is the standard deviation in ten random experiments with different training and testing data. Both the error and the distribution shift increase at a lower speed in $\mathrm{TR}$ simulation than in OLS simulations. This suggests that the better performance of $\mathrm{TR}$ algorithm in reaction-diffusion equations is partially caused by improving the distribution shift issue.
@@
We summarize the comparison results in Fig. 4. Here, solid lines represent the errors along the trajectories and dashed lines represent the distribution shift. The shaded part is the standard deviation over ten random experiments with different splits of the training and testing data. As can be observed, compared to our method the OLS baseline incurs greater errors and distribution shifts. Moreover, this difference increases with time. In contrast, consistent with our analysis in the simplified setting, our algorithm mitigates both the distribution shift and the trajectory error.
@@
Next, we consider the Navier-Stokes equation Eq. 3 as the test case of our algorithm Algorithm 1. In this case, the resolved model is an explicit scheme to evolve the velocity field while unresolved model is calculated via a fast CNN surrogate, which can accelerate the simulation for large grid systems. The simulation domain is rectangular with an aspect ratio $1/4$. On the upper and lower boundary we impose the no-slip boundary condition on the velocity and on the inlet, i.e. $x=0$ we specify the velocity field $\mathbf{u}$. Moreover, at the outlet, i.e. $x=n$ we pose zero-gradient conditions on both the horizontal velocity and pressure, i.e. $\frac{\partial p}{\partial n}=\frac{\partial u}{\partial n}=0$. The inflow boundary is defined according to following functional forms:
(50) $\mathbf{u}(0,y,t)=\begin{pmatrix}u(0,y,t)\\ v(0,y,y)\end{pmatrix}=\begin{pmatrix}\exp\{-50(y-y_{0})^{2}\}\\ \sin t\cdot\exp\{-50(y-y_{0})^{2}\}\end{pmatrix}$
@@
The training configuration is calculated using the projection method  [8] with a staggered grid, i.e. the pressure is placed at the center of each cell and horizontal (vertical) velocity is placed at vertical (horizontal) edge of each cell. The jet location $y_{0}$ is selected from $0.3$ to $0.7$ uniformly.
@@
Again, we compare the performance of our algorithm and the OLS method. As illustrated in Fig. 5, naively using OLS to estimate the unresolved models will generally lead to an error blow-up, whereas using our regularized estimator we can obtain a reasonable simulation along the whole time horizon. Moreover, the trend of the trajectory error fits that of the distribution shift. In particular, the error and distribution shift blow-ups occur simultaneously. Our algorithm again decreases the trajectory error by mitigating the distribution shift problem. Moreover, the blow-up of the error in the simulated dynamics is not explained in Theorem 3.4 as the theory is established for linear dynamics while finite-time blow-up is known to be a special feature of nonlinear dynamics.
We also notice the unphysical noise appearing in the numerical experiments of the Navier-Stokes equation, i.e. Fig. 5 b. which is also observed in [31, 32, 61, 43] when performing data-driven RANS modeling. While this is not the main focus of this paper which deals with the distribution shift issue, we observe that such noise mainly occurs in the simulation of the NS equation instead of reaction-diffusion equations. Therefore, one possible explanation is that this noise is caused by transport-dominated equations and diffusion can reduce this noise in diffusion-dominated equations.
@@
Figure 5: Simulation of the Navier-Stokes equation and configurations at time step: 0, 60, 120, 180, 240, (a) ground truth; (b) simulated fluid velocity field using OLS estimator; (c) simulated fluid velocity field using $\mathrm{TR}$ estimator; (d) the comparison of the OLS estimator and the regularized one. The solid lines represent the trajectory error and the dashed lines represent the distribution shift calculated usingusing function $F$ defined in Eq. 20. We observe that the OLS estimator leads to error (and DS) blow-up at around 270 steps, while our method remains stable. Again, the trends of error and distribution shift are highly correlated, and our method that controls distribution shift leads to error control, consistent with our analyses.
@@ 4.2 Performance under distribution shift of different magnitudes
We now demonstrate concretely that our method brings more significant improvements when the intrinsic distribution shift is more severe. Intuitively speaking, since our algorithm penalizes the deviation of the simulated trajectories to the data manifold while OLS ignores this, it would be reasonable to expect that our algorithm will outperform OLS more in problems where such deviations are large. A sanity check showing different distribution shift extents under different parameters is provided in Figure 3 of supplementary materials.
Here, we introduce two more baselines for ablation studies. Compared with the OLS algorithm, these algorithms both add some regularization to the loss function while their regularizations are general and not dynamics-specific compared to our regularization. Consequently, any improved performance in our algorithm over these regularized algorithms suggest that targeted regularization, which takes into account the structure of the resolved part, effectively balances accuracy and stability.
The first baseline regularizes the original least squares objective by a term that quantifies some complexity of the network model, e.g.
(51) $\phi_{\text{mOLS}}=\arg\min\mathbb{E}_{\mathbf{u}\sim\rho_{0}}\left\|\phi_{\theta}(\mathbf{u})-\phi(\mathbf{u})\right\|_{2}^{2}+\lambda\left\|\theta\right\|_{2}^{2}.$
Here the naive $L^{2}$ regularized algorithm is implemented by adjusting the weight decay [34] parameter of the Adam solver during the training process. The corresponding numerical results are labeled by “mOLS” (“m” for modified).
@@
(52) $\phi_{\text{aOLS}}=\arg\min\mathbb{E}_{\mathbf{u}\sim\rho_{\xi}}\left\|\phi_{\theta}(\mathbf{u})-\phi(\mathbf{u})\right\|_{2}^{2}.$
Notice that the training data follows the distribution $\rho_{\epsilon}=\rho_{0}*\xi$ where * denotes the convolution operator and $\xi$ is some random noise which we take to be Gaussian random variables with an adjustable variance. In our implementation, we generate the noise and add it to the input during each epoch, so the inputs are not the same across epochs. Moreover, this noise $\xi$ should be distinguished from the observational noise $\epsilon$ we introduced in Section 3. This method has been used in several previous works  [54, 40] to improve stability and prevent the error from accumulating too fast. As this technique is similar to adversarial attack and data augmentation [52, 14, 13], we name this estimator and algorithm as “aOLS” (“a” stands for adversarial). We slightly abuse the terminology here by naming this random perturbation as some kind of adversarial attack.
Table 1: Performance on reaction-diffusion equations over 10 experiments.
@@
In Table 1, we test the algorithms under different diffusion coefficient $\gamma=0.05$, $0.10,0.15,0.20,0.25$. We calculate the relative trajectory error at time $T=1000$, i.e. $\frac{\left\|\mathbf{u}_{k}-\widehat{\mathbf{u}}_{k}\right\|_{2}}{\left\|\mathbf{u}_{k}\right\|_{2}}$ over ten random splits of the training and test data. The last row calculates the performance difference between our method and the OLS baseline. We observe that under almost all the settings our method out-performs OLS, mOLS, aOLS baselines, and comparing horizontally, we see that the relative improvement decreases with $\gamma$, as the diffusion term mitigates the distribution shift.
We now consider the Navier-Stokes equation. Since the baseline method quickly leads to error blow-up, we use another comparison criterion. We define a stopping time $t_{K}=\arg\max_{t}\frac{\left\|\mathbf{u}_{t}-\widehat{\mathbf{u}}_{t}\right\|_{2}}{\left\|\mathbf{u}_{t}\right\|_{2}}\leq K$ where $K$ is an error threshold to be determined during the experiments. We calculate the first time the trajectory error reaches a threshold under different Reynolds numbers and different mesh sizes. The results are shown in Table 2. Under almost all the scenarios, our method outperforms all the baseline algorithms by a larger $t_{K}$. This validates the effectiveness of our algorithm. Comparing Table 2 horizontally, one finds that the improvement of our method is also increasing concerning the Reynolds number. As the flow field generally becomes more complex (and sensitive) with increasing Reynolds number, our method is expected to bring bigger improvements. This is consistent with our experiments.
@@
$\mathrm{TR}$ $\mathbf{506\pm 18}$ $\mathbf{497\pm 28}$ $\mathbf{482\pm 29}$ $\mathbf{442\pm 21}$ $\mathbf{452\pm 13}$
@@
$\mathrm{TR}$ $636\pm 35$ $\mathbf{602\pm 28}$ $585\pm 23$ $\mathbf{422\pm 35}$ $\mathbf{402\pm 28}$